# AHSG: Adversarial Attacks on High-level Semantics in Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have garnered significant interest among researchers due to their impressive performance in graph learning tasks. However, like other deep neural networks, GNNs are also vulnerable to adversarial attacks. In existing adversarial attack methods for GNNs, the metric between the attacked graph and the original graph is usually the attack budget or a measure of global graph properties. However, we have found that it is possible to generate attack graphs that disrupt the primary semantics even within these constraints. To address this problem, we propose a Adversarial Attacks on High-level Semantics in Graph Neural Networks (AHSG), which is a graph structure attack model that ensures the retention of primary semantics. The latent representations of each node can extract rich semantic information by applying convolutional operations on graph data. These representations contain both task-relevant primary semantic information and task-irrelevant secondary semantic information. The latent representations of same-class nodes with the same primary semantics can fulfill the objective of modifying secondary semantics while preserving the primary semantics. Finally, the latent representations with attack effects is mapped to an attack graph using Projected Gradient Descent (PGD) algorithm. By attacking graph deep learning models with some advanced defense strategies, we validate that AHSG has superior attack effectiveness compared to other attack methods. Additionally, we employ Contextual Stochastic Block Models (CSBMs) as a proxy for the primary semantics to detect the attacked graph, confirming that AHSG almost does not disrupt the original primary semantics of the graph.
Introduction
Graph structures are used to model many real-world relationships, such as molecular structures (Dai et al. 2019), social relationships (Bian et al. 2020), network flows (Shen et al. 2022), and transportation (Zhou et al. 2023). Learning effective representations of graphs and applying them to solve downstream tasks has become increasingly important. In recent years, GNNs (Kipf and Welling 2017; Veličković et al. 2018; Hamilton, Ying, and Leskovec 2017)have achieved significant success in graph representation learning. GNNs follow the message-passing mechanism (Gilmer et al. 2017), where node embeddings are obtained by aggregating and transforming the embeddings of their neighbours. Due to their excellent performance, GNNs have been applied to various analytical tasks, including node classification (Jin et al. 2021a; Wu, Pan, and Zhu 2021), link prediction (Xiong et al. 2022; Ma et al. 2020), and graph classification (Gao et al. 2021; Errica et al. 2022).
@@ Introduction
GNNs have achieved significant breakthroughs in many graph-related tasks. However, recent studies have shown that similar to traditional deep neural networks, GNNs suffer from poor robustness when facing specially designed adversarial attacks. Attackers can generate graph adversarial perturbations to deceive GNNs by manipulating graph structures and node features (Dai et al. 2018; Xu et al. 2019; Zügner and Günnemann 2019; Zügner, Akbarnejad, and Günnemann 2018; Liu et al. 2022; Zhang et al. 2022), or generating new nodes and adding them to the original graph (Fang et al. 2024; Chen et al. 2022). By understanding the ways in which models are vulnerable to attacks, researchers can design more robust models (Zhu et al. 2019; Wu et al. 2019; Jin et al. 2020, 2021b; Entezari et al. 2020) that perform better when confronted with adversarial samples.
A core principle of attacking neural networks is that the attacker should preserve the primary semantics of the original data after adding perturbations. In other words, human observers should perceive the samples before and after perturbation as semantically equivalent. For example, an image should still be recognized by humans as the same object even after perturbation. This semantic invariance ensures the practical significance of the attack. To achieve semantic invariance in adversarial examples, adversarial attacks usually require the attacker to make only minimal perturbations. Therefore, most existing research on graph adversarial attacks (Dai et al. 2018; Xu et al. 2019; Zügner and Günnemann 2019; Zügner, Akbarnejad, and Günnemann 2018; Liu et al. 2022) restricts the attacker to modify only a limited number of edges or nodes. This limitation is known as the attack budget. These studies assume that within a limited attack budget, the semantics of the adversarial example do not change. However, it is questionable whether attack models of the kind are able to preserve the primary semantics. For example, the proportion of low-degree nodes is large in real-world graphs, and small attack budget still allows the original neighbours of low-degree nodes to be completely deleted. Once the edges of low-degree nodes are destroyed, the semantics of these nodes are likely to be destroyed. It is not enough to limit the change of semantics only by using the attack budget.
@@
To further constrain semantic changes, a few studies have introduced metrics beyond attack budget limitations. (Zügner, Akbarnejad, and Günnemann 2018; Chen et al. 2022) suggest using different global graph properties as proxies for semantics, such as degree distribution (Zügner, Akbarnejad, and Günnemann 2018) and homophily (Chen et al. 2022). However, we find that even under such constraints, the primary semantics can still be disrupted. Figure 1 illustrates this phenomenon with a binary classification task. Assume we have an undirected graph. Select any two edges $e_{1}=(i,j)$ and $e_{2}=(u,v)$ from the edge set of the graph. Now, we replace the two edges with $\widehat{e_{1}}=(i,v)$ and $\widehat{e_{2}}=(u,j)$. Since $\deg(i)$, $\deg(j)$, $\deg(u)$, and $\deg(v)$ remain unchanged, this process preserves the degree distribution of the graph. We can continue to modify the edges according to Figure 1 until the attack budget is exhausted. We successfully modifies the semantics of nodes 3 and 8 in Figure 1 while maintaining the overall graph distribution, demonstrating the infeasibility of method (Zügner, Akbarnejad, and Günnemann 2018). Therefore, neither the attack budget nor the additional constraints on graph properties can effectively preserve the primary semantics of the graph. Current research lacks the fundamental constraint of maintaining primary semantics. To address this issue, we propose a high-level semantic adversarial attack targeting graph neural networks. This method aims to generate attack graphs that preserve primary semantics within limited attack budget.
@@
The semantics of the graph is contained in the latent representations of graph neural network, so we control the change of the latent representations to control the semantic change of the attacked graph. Existing attack methods (Dai et al. 2018; Xu et al. 2019; Zügner and Günnemann 2019; Zügner, Akbarnejad, and Günnemann 2018; Liu et al. 2022; Zhang et al. 2022) often treat graph neural networks as black boxes, focusing only on the input and output, while ignoring the rich semantic information contained in the hidden layer (its own node features as well as the structural information of the surrounding neighborhood). While many of them represent the primary semantics of an object in a fixed task, some of them are secondary. For example, in social networks, the height of person nodes is secondary to the task of predicting friend relationships, while the interests of person nodes are primary. These secondary semantics may play an inappropriately important role in the model predictions. In an ideal situation, the construction of adversarial examples should lead the model to misjudge through exploring and exploiting the secondary semantics while keeping the primary semantics unchanged.
@@
We endeavor to seek the attacked graph which preserves the primary semantics from the latent representations that contain semantics. Subsequently, how can we acquire the latent representations where primary semantics remain unchanged while secondary semantics change? Nodes of the same class have similar primary semantics in a fixed task, so a linear combination of representations of them can still retain their primary semantics. We use the latent representations after linear combination as input to the lower network, then use gradient ascent to find latent representations with attack effects. To generate adversarial samples, a first-order optimization algorithm is employed to map the perturbed latent representations with attack effects to the attacked graph. Finally, extensive experiments validate the attack performance of AHSG on different datasets and against different defense methods.
@@
  • To overcome the challenge of preserving primary semantics in attacked graph, we propose a new strategy based on a novel perspective of latent semantic representations in graph neural network, which cleverly utilizes the same-class latent representations with similar primary semantics as the perturbation range.
  • In the setting of evasion attacks, we introduce AHSG. By extracting continuous latent representations from discrete graph data through convolutional operations, we perform a linear combination of same-class node representations as perturbations. Then, we use a first-order optimization algorithm to reconstruct adversarial samples from the perturbed representations.
  • Through comprehensive experiments on multiple acknowledged benchmark datasets, we compare AHSG with current advanced attack models and use various defense models for evaluation. The results demonstrate that AHSG achieves excellent attack performance under different attack budgets. Furthermore, semantic detection experiments confirm that AHSG preserves the primary semantics of graph data in fixed tasks.
Preliminary
@@
GCN
Graph Convolutional Network (GCN) (Kipf and Welling 2017) is one of the classic models based on graph learning and has been widely applied to various tasks. The principle of GCN involves a message passing mechanism, where each node updates its representation by aggregating information from itself and its neighboring nodes using pre-defined deterministic rules (e.g., Laplacian matrix, adjacency matrix, and attention). The process of message passing can be represented as follows:
$H^{(l+1)}=\sigma(\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}H^{(l)}W_{(l)}).$ (1)
Here, the node representation $H^{(l)}$ is updated to $H^{(l+1)}$ after message passing. In Equation (1), $\widetilde{A}=A+I_{N}$ and $\widetilde{D}_{ii}=\sum_{j}\widetilde{A}_{ij}$, where $\widetilde{D}_{ii}$ is equal to the degree of node $v_{i}$ plus 1. $\sigma$(.) is a non-linear activation function, such as ReLU. The initial node representation $H^{(0)}$ is set to the node features $X$. $W_{(l)}$ $(l=1,2,\cdots,L)$ are the weight matrices. Typically, a fully connected layer with a softmax function is used for classification after $L$ graph convolutions. Generally, a two-layer GCN is considered for node classification tasks. Thus, the model can be described as:
$Z=f(G(X,A))=\text{softmax}(\hat{A}\sigma(\hat{A}XW_{(1)})W_{(2)}),$ (2)
where $\hat{A}=\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$ is the symmetrically normalized adjacency matrix.
Graph Adversarial Attack
@@
$\max_{\hat{G}\in\mathrm{\Phi}(G)}L\left(f_{\theta^{\ast}}(\hat{G}(\hat{A},\hat{X})),Y\right)$ (3)
@@
Figure 2: The framework of AHSG.We use the same color scheme to represent latent representations of same-class nodes. The color indicates primary semantics, while color depth represents secondary semantics.
Methodology
@@
The hidden layers of an unperturbed surrogate GCN encapsulate both the information of the node itself and that of its surrounding neighborhood after training. Therefore, we consider the output from the first convolutional layer and activation function as the semantics of each node. This process can be formally described as follows:
$H_{1}=\sigma\left(n(A)XW_{(1)}^{\ast}\right)$ (4)
@@
where $f_{W_{1,2}}\left(G\left(A,X\right)\right)$ is a shorthand for the expression in Equation (2), $L\left(.,.\right)$ is the loss function, commonly the cross-entropy loss. $n(A)=\widetilde{D}^{-\frac{1}{2}}(A+I)\widetilde{D}^{-\frac{1}{2}}$.
@@
(Wang et al. 2021) addresses the challenge of imbalanced datasets by extracting the feature shifts of frequent-class entities and applying them to rare-class entities. In this approach, the latent representation is assumed to be $h=P+T$, where $P$ denotes the data distribution center (i.e., the similar primary semantics of same-class entities under the fixed task), and $T$ represents the feature shift (i.e., the secondary semantics of the entities in the task). Inspired by this, we make the following assumption about the composition of the latent representation in our work:
@@
where $h_{j}^{i}$ denotes the latent representation vector of the $j$-th node in class $i$. It is composed of two parts: the primary semantics $eP(i)$, which is crucial for the classification task, and the secondary semantics $fT(j)$, which is less decisive. All nodes of the same class share the same primary semantic base $P(i)$, but may have different secondary semantic bases $T(j)$. Since each class contains some nodes, there exists a set where the primary semantics are preserved while the secondary semantics change. When perturbing the latent representation $h_{j}^{i}$ of each node, we assign a weight $\alpha_{k}$ to each representation of the same class. Summing these weighted representations yields the perturbed representation $\hat{h}_{j}^{i}$.
@@
From Equation (5), the representation $\hat{h}_{j}^{i}$ can be derived as follows:
$\hat{h}_{j}^{i}$ $=\frac{\sum_{k=1}^{j}\alpha_{k}h_{k}^{i}}{\sum_{k=1}^{j}\alpha_{k}}$ (7)
$=\frac{\sum_{k=1}^{j}\alpha_{k}\left(d_{k}P(i)+e_{k}T(k)\right)}{\sum_{k=1}^{j}\alpha_{k}}$
@@
$\text{loss}_{1}=-L(Y,\text{softmax}(n(A)\widehat{H_{1}}(\alpha)W_{(2)}^{\ast}))$ (8)
$-\beta\,\text{sim}(H_{1},\widehat{H_{1}}(\alpha)).$
$\widehat{H_{1}}(\alpha)=\left[\hat{h}_{j}^{i}\right]=\left[\frac{\sum_{k=1}^{j}\alpha_{k}h_{k}^{i}}{\sum_{k=1}^{j}\alpha_{k}}\right].$ (9)
To use a smaller attack budget when reconstructing adversarial samples with perturbed representations, we introduce a penalty term $\text{sim}(H_{1},\widehat{H_{1}}(\alpha))$. This term controls the similarity between the perturbed hidden layer and the original hidden layer, thereby limiting the perturbations to a certain range. Notably, it also prevents the coefficient of the primary semantics base from becoming too small. $\beta>0$ is a hyperparameter that controls the strength of the penalty.
@@
To construct an adversarial sample, we need to map the perturbed latent representation to the graph. Let $\hat{A}$ denote the attacked graph. We introduce a boolean symmetric matrix $S\in\{0,1\}^{N\times N}$ to encode whether an edge in $G$ is modified. Specifically, an edge $(i,j)$ is modified (added or removed) if $S_{ij}=S_{ji}=1$. Otherwise, if $S_{ij}=S_{ji}=0$, the edge $(i,j)$ remains undisturbed. Given the adjacency matrix $A$, its complement matrix $\bar{A}$ is defined as $\bar{A}=11^{T}-I-A$, where $I$ is the identity matrix, and $1$ is a column vector of all ones. The term $(11^{T}-I)$ corresponds to a fully connected graph. Using the edge perturbation matrix $S$ and the complement matrix $\bar{A}$, Equation (10) provides the perturbed topology $\hat{A}$ of the graph $A$.
$\hat{A}=A+C\odot S,\quad C=\bar{A}-A.$ (10)
@@
$\text{loss}_{2}=-\text{sim}\left(\sigma(n(\hat{A})XW_{(1)}^{\ast}),\widehat{H_{1}}\right)\text{ s.t. }s\in S.$ (11)
[…]
Model Optimization

### v2

# AHSG: Adversarial Attack on High-level Semantics in Graph Neural Networks

@@ Abstract
Adversarial attacks on Graph Neural Networks aim to perturb the performance of the learner by carefully modifying the graph topology and node attributes. Existing methods achieve attack stealthiness by constraining the modification budget and differences in graph properties. However, these methods typically disrupt task-relevant primary semantics directly, which results in low defensibility and detectability of the attack. In this paper, we propose an Adversarial Attack on High-level Semantics for Graph Neural Networks (AHSG), which is a graph structure attack model that ensures the retention of primary semantics. By combining latent representations with shared primary semantics, our model retains detectable attributes and relational patterns of the original graph while leveraging more subtle changes to carry out the attack. Then we use the Projected Gradient Descent algorithm to map the latent representations with attack effects to the adversarial graph. Through experiments on robust graph deep learning models equipped with defense strategies, we demonstrate that AHSG outperforms other state-of-the-art methods in attack effectiveness. Additionally, using Contextual Stochastic Block Models to detect the attacked graph further validates that our method preserves the primary semantics of the graph.
1 Introduction
@@
GNNs Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2017)have achieved significant success in graph representation learning. Due to their excellent performance, GNNs have been applied to various analytical tasks, including node classification Jin et al. (2021a); Wu et al. (2021b), link prediction Xiong et al. (2022), and graph classification Gao et al. (2021). Recent studies have shown that similar to traditional deep neural networks, GNNs suffer from poor robustness when facing specially designed adversarial attacks. Attackers can generate graph adversarial perturbations to deceive GNNs by manipulating graph structures and node features Dai et al. (2018); Xu et al. (2019); Zügner and Günnemann (2019); Zügner et al. (2018); Liu et al. (2022); Zhang et al. (2022), or generating new nodes and adding them to the original graph Fang et al. (2024); Chen et al. (2022). By understanding the ways in which models are vulnerable to attacks, a range of defence methods are designed that perform better when confronted with adversarial samples Zhu et al. (2019); Wu et al. (2019); Jin et al. (2020, 2021b); Entezari et al. (2020).
@@
A core principle of attacking neural networks is that the attacker should preserve the primary semantics of the original data after adding perturbations. This semantic invariance ensures the practical significance of the attack. To achieve this, adversarial attacks usually require the attacker to make only minimal perturbations. Therefore, most existing research on graph adversarial attacks Dai et al. (2018); Xu et al. (2019); Zügner and Günnemann (2019); Zügner et al. (2018); Liu et al. (2022) restricts attackers to modifying a limited number of edges or nodes. However, it remains uncertain whether such attack models can effectively preserve the primary semantics. For example, real-world graphs typically have many low-degree nodes, and a small attack budget can completely remove their original neighbors. Once the edges of low-degree nodes are disrupted, their semantics are often lost. Therefore, preserving the primary semantics requires more than just managing the attack budget.
@@
To further constrain semantic changes, existing methods introduce metrics beyond attack budget limitations. Zügner et al. (2018); Chen et al. (2022) suggest using different global graph properties as proxies for semantics, such as degree distribution and homophily. However, even under such constraints, the primary semantics can still be disrupted. Figure 1 illustrates this phenomenon with a binary classification task. Assume we select two edges $e_{1}=(i,j)$ and $e_{2}=(u,v)$ from the edge set of an undirected graph, and replace them with $\widehat{e_{1}}=(i,v)$ and $\widehat{e_{2}}=(u,j)$. Since $\deg(i)$, $\deg(j)$, $\deg(u)$, and $\deg(v)$ remain unchanged, this process preserves the degree distribution of the graph. We can continue to modify the edges according to this rule until the attack budget is exhausted. Finally, we successfully modifies the semantics of nodes 3 and 8 while maintaining the overall graph distribution, demonstrating the infeasibility of method Zügner et al. (2018). Therefore, neither the attack budget nor the additional constraints on graph properties can effectively preserve the primary semantics of the graph. Current research lacks a fundamental constraint aimed at maintaining primary semantics.
@@
To address the above question, we propose a method that generates attack graph preserving primary semantics within limited attack budget. First, we need to find tools that represent semantics. Existing attack methods Dai et al. (2018); Xu et al. (2019); Zügner and Günnemann (2019); Zügner et al. (2018); Liu et al. (2022); Zhang et al. (2022) often treat graph neural networks as black boxes, focusing only on the input and output, while ignoring the rich semantic information contained in the hidden layers. The semantics of the graph are considered to be contained in the latent representations Wu et al. (2021a), so we control the change of the latent representations to control the semantic change of the attacked graph. Although the hidden layer representations in common tasks often encapsulate the primary semantics of the analyzed objects, task-irrelevant secondary semantics are not completely filtered out. These secondary semantics may play an inappropriately role in the model predictions. In this context, the construction of adversarial examples should lead the model to misjudge through exploring and exploiting the secondary semantics while keeping the primary semantics unchanged.
@@
Subsequently, how can we get latent representations where primary semantics remain unchanged while secondary semantics change? Inspired by Wang et al. (2021), the entities within the same category share similar primary semantics while exhibiting diverse secondary semantics. Therefore, their linearly combined representations can preserve the majority of the consensus primary semantics, while the deviations in secondary semantics can serve as the basis for crafting attacks. We use the latent representations after linear combination as input to the deeper network layers, then use gradient ascent to find latent representations with attack effects. To generate adversarial samples, a first-order optimization algorithm is employed to map the perturbed latent representations with attack effects to the attacked graph. Finally, this secondary-semantics-targeted attack method achieves lower defensibility and detectability.
@@
  • We propose a novel adversarial attack framework AHSG based on the graph semantics, which utilizes the class-shared latent representations with similar primary semantics as the perturbation tools.
  • We introduce a technique for reconstructing adversarial samples from perturbed representations by extracting deviations in secondary semantics from the linear combination of class-shared representations, while preserving the primary semantics.
  • We validate AHSG on multiple acknowledged benchmark datasets, demonstrating substantial improvements under different attack settings. Semantic detection experiments confirm that AHSG preserves the primary semantics of graph data.
2 Related Work
2.1 Adversarial Attack on GNNs
In graph attacks, Nettack Zügner et al. (2018) introduces the first adversarial attack method on graphs. Meta-Self Zügner and Günnemann (2019) treats the input graph as a hyperparameter to be learned and modifies one edge per iteration. PGD Xu et al. (2019) overcomes the challenge of attacking discrete graph structure data. Dai et al. (2018) employ reinforcement learning, gradient-based greedy algorithms, and genetic algorithms to attack GNNs in various scenarios. EpoAtk Lin et al. (2020) improves the effectiveness of the attack by bypassing potential misinformation from the maximal gradient. GraD Liu et al. (2022) generates unweighted gradients on the graph structure, unaffected by node confidence, fully utilizing the attack budget. However, the issue of altering graph semantics during an attack remains underexplored. Although the attack budget is minimal relative to the entire graph, it can still lead to significant changes in the semantics of the attacked graph. Some works, such as Nettack Zügner et al. (2018) and HAO Chen et al. (2022), maintain certain graph properties to achieve imperceptibility. However, these constraints are necessary but not sufficient to preserve semantics.
2.2 Adversarial Defense on GNNs
Some studies develop corresponding defense strategies based on the characteristics of the attacks. Jaccard Wu et al. (2019) calculates the Jaccard similarity of connected node pairs and retains only those links with high similarity. Svd Entezari et al. (2020) utilizes low-order approximations of the graph to enhance the performance of GCN against adversarial attacks. ProGNN Jin et al. (2020) jointly learns the graph structure and a robust GNN model from the perturbed graph. The SimPGCN Jin et al. (2021b) framework enhances GCN robustness by effectively preserving node similarity. RGCN Zhu et al. (2019) proposes that latent representations based on Gaussian distributions can effectively absorb the impact of adversarial attacks. In the experiment, we use AHSG along with baseline methods to attack the aforementioned defense models.
3 Preliminary
@@
3.1 Graph Neural Networks
GNNs are specifically designed to handle graph-structured data. They leverage the message passing mechanism to aggregate information from nodes and their neighbors, thereby learning feature representations for nodes, edges, and the graph as a whole. The formula for message passing can be described as follows:
$h_{v}^{(l)}=\text{UD}^{(l)}\left(h_{v}^{(l-1)},h_{t}\right).$ (1)
The representation of node $v$ at the $l$-th layer, denoted as $h_{v}^{(l)}$, is computed by applying an update function $\text{UD}^{(l)}$ to its previous representation $h_{v}^{(l-1)}$ and the aggregated information $h_{t}$ from its neighbors.
$h_{t}=\text{AG}^{(l)}\left(\left\{\text{MG}^{(l)}\left(h_{u}^{(l-1)},h_{v}^{(l-1)}\right)\mid u\in\mathcal{N}(v)\right\}\right).$ (2)
The aggregation function $\text{AG}^{(l)}$ collects messages generated by the message function $\text{MG}^{(l)}$, which computes the message from each neighboring node $u$ to node $v$ based on their respective representations $h_{u}^{(l-1)}$ and $h_{v}^{(l-1)}$. Here, $\mathcal{N}(v)$ represents the set of neighboring nodes of $v$. The initial node representation $h^{(0)}$ is set to the node features $X$. The learned representations are used as inputs for downstream tasks.
3.2 Graph Adversarial Attack
@@
$\max_{\hat{G}\in\mathrm{\Phi}(G)}\mathcal{L}\left(f_{\theta^{\ast}}(\hat{G}(\hat{A},\hat{X})),Y\right)$ (3)
@@
Figure 2: The framework of AHSG. It consists of three steps: extracting semantics from graph, constructing adversarial latent representations, and mapping adversarial latent representation to attacked graph.
4 Methodology
@@ 4.1 Architecture of AHSG
The hidden layers of an unperturbed surrogate GNN encapsulate both the information of the node itself and that of its surrounding neighborhood after training. Therefore, we select the output of the $l$ layer as the semantics of each node. This process can be formally described as follows:
$H_{l}=GNN_{\{1\sim l\}}\left(A,X,W_{(1,\cdots,l)}^{\ast}\right)$ (4)
@@
where $\mathcal{L}\left(.,.\right)$ is the loss function, commonly the cross-entropy loss. $W_{(1,\cdots,L)}$ are weight matrices. $GNN_{\{1\sim l\}}$ refer to the first $l$ layers of the GNN.
@@
Wang et al. (2021) address the challenge of imbalanced datasets by extracting the feature shifts of frequent-class entities and applying them to rare-class entities. In this approach, the latent representation is assumed to be $h=P+T$, where $P$ denotes the data distribution center (i.e., the similar primary semantics of same-class entities under the fixed task), and $T$ represents the feature shift (i.e., the secondary semantics of the entities in the task). Inspired by this, we make the following assumption about the composition of the latent representation in our work:
@@
where $h_{j}^{i}$ denotes the latent representation vector of the $j$-th node in class $i$. It is composed of two parts: the primary semantics $dP(i)$, which is crucial for the classification task, and the secondary semantics $eT(j)$, which is less decisive. All nodes of the same class share the same primary semantic base $P(i)$, but may have different secondary semantic bases $T(j)$. Since each class contains some nodes, there exists a set where the primary semantics are preserved while the secondary semantics change. When perturbing the latent representation $h_{j}^{i}$ of each node, we assign a weight $\alpha_{k}$ to each representation of the same class. Summing these weighted representations yields the perturbed representation $\hat{h}_{j}^{i}$.
@@
Using Equation (5), the representation $\hat{h}_{j}^{i}$ in Equation (6) can be derived as follows:
$\hat{h}_{j}^{i}$ $=\frac{\sum_{k=1}^{j}\alpha_{k}h_{k}^{i}}{\sum_{k=1}^{j}\alpha_{k}}=\frac{\sum_{k=1}^{j}\alpha_{k}\left(d_{k}P(i)+e_{k}T(k)\right)}{\sum_{k=1}^{j}\alpha_{k}}$ (7)
@@
$\mathcal{L}_{1}=-\mathcal{L}(Y,GNN_{\{l\sim L\}}\left(A,\widehat{H_{l}}(\alpha),W_{(l,\cdots,L)}^{\ast}\right))$ (8)
$-\beta\,\text{sim}(H_{l},\widehat{H_{l}}(\alpha)),$
[…]
$\widehat{H_{l}}(\alpha)=\left[\hat{h}_{j}^{i}\right]=\left[\frac{\sum_{k=1}^{j}\alpha_{k}h_{k}^{i}}{\sum_{k=1}^{j}\alpha_{k}}\right].$ (9)
To use a smaller attack budget when reconstructing adversarial samples with perturbed representations, we introduce a penalty term $\text{sim}(H_{l},\widehat{H_{l}}(\alpha))$. This term controls the similarity between the perturbed hidden layer and the original hidden layer, thereby limiting the perturbations to a certain range. Notably, it also prevents the coefficient of the primary semantics base from becoming too small. $\beta>0$ is a hyperparameter that controls the strength of the penalty.
@@
To construct an adversarial sample, we need to map the perturbed latent representation to the graph. Let $\hat{A}$ denote the attacked graph. We introduce a binary perturbation matrix $S\in\{0,1\}^{N\times N}$ to encode whether an edge in $G$ is modified. Specifically, an edge $(i,j)$ is modified (added or removed) if $S_{ij}=S_{ji}=1$. Otherwise, if $S_{ij}=S_{ji}=0$, the edge $(i,j)$ remains undisturbed. Given the adjacency matrix $A$, its complement matrix $\bar{A}$ is defined as $\bar{A}=11^{T}-I-A$, where $I$ is the identity matrix, and $1$ is a column vector of all ones. The term $(11^{T}-I)$ corresponds to a fully connected graph. Using the edge perturbation matrix $S$ and the complement matrix $\bar{A}$, Equation (10) provides the perturbed topology $\hat{A}$ of the graph $A$.
$\hat{A}=A+C\odot S,\quad C=\bar{A}-A,$ (10)
where $\odot$ denotes the element-wise product. In the above expression, the positive entries of $C$ indicate edges that can be added to the graph $A$, while the negative entries of $C$ indicate edges that can be removed from $A$. Due to the difficulty in solving the problem under the binary constraint $\{0,1\}$, this constraint is relaxed to the continuous interval $[0,1]$. Based on Equation (4), the latent representation is computed by using the modified graph structure $\hat{A}$, which is similar to $\widehat{H_{l}}$. Hence, we give the mapping loss as follows:
$\mathcal{L}_{2}=-\text{sim}\left(GNN_{\{1\sim l\}}\left(\hat{A},X,W_{(1,\cdots,l)}^{\ast}\right),\widehat{H_{l}}\right)\text{ s.t. }s\in S,$ (11)
@@
4.2 Model Optimization